# “Sometimes You Just Gotta Risk It for the Biscuit [Uncaptioned image]”: A Portrait of Student Risk-Taking

### v2

@@ 6. Conclusions
We conducted an experiment looking at the risk-taking behavior of students in an introductory programming course inspired by an earlier study by Graf-Vlachy (Graf-Vlachy, 2023) that looked at the risk-taking behavior of professional software engineers. We were able to replicate earlier findings suggesting that the framing of choices correlates with risk-taking, and that framing the choices as “losses” as opposed to “gains” correlates with making riskier choices – choices where the outcome is based on chance as opposed to a guaranteed outcome. Another interesting finding was that students were in general a lot more prone to taking risks than professional software engineers. Interestingly, students who performed better in the course, and who enjoyed programming more, were somewhat more likely to make risky choices, while whether students found programming difficult or not did not correlate statistically significantly with risk-taking. Lastly, we found that students justified their risk-related choices in different ways, such as based on their risk tolerance, strategic calculations, and emotional factors. Altogether, our study sheds light on risk-taking behavior of students in computing, which has been underexplored in previous work.

### v3

@@ 6. Conclusions
We conducted an experiment looking at the risk-taking behavior of students in an introductory programming course inspired by an earlier study by Graf-Vlachy (Graf-Vlachy, 2023) that looked at the risk-taking behavior of professional software engineers. We were able to replicate earlier findings suggesting that the framing of choices correlates with risk-taking, and that framing the choices as “losses” as opposed to “gains” correlates with making riskier choices – choices where the outcome is based on chance as opposed to a guaranteed outcome. Another interesting finding was that students were in general more inclined to take risks than professional software engineers. Interestingly, students who performed better in the course, and who enjoyed programming more, were somewhat more likely to make risky choices, while whether students found programming difficult or not did not correlate statistically significantly with risk-taking. Lastly, we found that students justified their risk-related choices in different ways, such as based on their risk tolerance, strategic calculations, and emotional factors. Altogether, our study sheds light on risk-taking behavior of students in computing, which has been underexplored in previous work.